OpenStudy (anonymous):

I need help with these questions please. FInd the LCM of x^2 + 10x , x^2 - 10x I will draw other ones below

jimthompson5910 (jim_thompson5910):

factor x^2 + 10x to get ______

OpenStudy (anonymous):

x(x + 10)

jimthompson5910 (jim_thompson5910):

factor x^2 - 10x to get _____

OpenStudy (anonymous):

x(x - 10)

jimthompson5910 (jim_thompson5910):

the unique factors are: x, x+10, x-10

jimthompson5910 (jim_thompson5910):

multiply all of these unique factors to get x(x+10)(x-10)

jimthompson5910 (jim_thompson5910):

and that's the LCM (factored out)

jimthompson5910 (jim_thompson5910):

you can choose to expand it out, but you're pretty much done

OpenStudy (anonymous):

wow thank you! Could you help me with two others?

jimthompson5910 (jim_thompson5910):

sure

jimthompson5910 (jim_thompson5910):

I can help

OpenStudy (anonymous):

\[Solve R = \frac{ gy }{ g + y } for y \]

jimthompson5910 (jim_thompson5910):

\[\Large R = \frac{ gy }{ g + y }\] \[\Large R(g+y) = gy\] \[\Large Rg+Ry = gy\] \[\Large Rg = gy-Ry\] \[\Large Rg = (g-R)y\] See what to do from here?

OpenStudy (anonymous):

I need to get y alone for I would divide both sides by (g - R)??

jimthompson5910 (jim_thompson5910):

exactly, so you would get \[\Large R = \frac{ gy }{ g + y }\] \[\Large R(g+y) = gy\] \[\Large Rg+Ry = gy\] \[\Large Rg = gy-Ry\] \[\Large Rg = (g-R)y\] \[\Large \frac{Rg}{g-R} = y\] \[\Large y = \frac{Rg}{g-R}\]

OpenStudy (anonymous):

\[Simplify the expression leave \in factored form. \frac{ 2y }{ y(2y - 1) } + \frac{ 1 }{ 2y - 1 }\]

jimthompson5910 (jim_thompson5910):

I'll get you started

OpenStudy (anonymous):

Thanks

jimthompson5910 (jim_thompson5910):

the LCD is y(2y - 1), so you need to get each denominator equal to the LCD the first denominator is already there, so no need to worry about the first fraction the second denominator is not equal to the LCD, so you need to multiply top and bottom by 'y' to get to the LCD

jimthompson5910 (jim_thompson5910):

so you'll go from \[\Large \frac{ 2y }{ y(2y - 1) } + \frac{ 1 }{ 2y - 1 }\] to \[\Large \frac{ 2y }{ y(2y - 1) } + \frac{ y }{ y(2y - 1) }\]

jimthompson5910 (jim_thompson5910):

the denominators are now equal, which means you can add the two fractions

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

\[\frac{ 3y }{ y(2y - 1) } ?\]

jimthompson5910 (jim_thompson5910):

now the y terms will cancel leaving you with \[\Large \frac{ 3 }{ 2y - 1}\]

OpenStudy (anonymous):

Cool.. Thank you for your help!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

If I ever have another question could I ask you? I like how you help but also let me work on the problem

jimthompson5910 (jim_thompson5910):

sure I'd love to help you out if you have another question

OpenStudy (anonymous):

\[Solve and Check \frac{ 3 }{ 1 - x } = \frac{ 4 }{ 1 + x }\]

OpenStudy (anonymous):

I rewrote it as 3(1 + x) = 4(1 - x) but than I am not sure where to go from there

jimthompson5910 (jim_thompson5910):

You want to isolate x, so you have to get all the x terms to one side (and everything else to the other) 3(1 + x) = 4(1 - x) 3 + 3x = 4 - 4x ... distribute 3+3x + 4x = 4 3x+4x = 4 - 3 what's next?

OpenStudy (anonymous):

Wouldnt it be 3 + 3x = 4 - 4x 3 + 3x + 4x = 4 3 + 7x = 4 -3 -3 7x = 1 x = 1/7

jimthompson5910 (jim_thompson5910):

good 3x+4x = 4-3 7x = 1 x = 1/7

OpenStudy (anonymous):

Thank you! That is it for now. I will be back again next Sunday! Hope to see you here.

jimthompson5910 (jim_thompson5910):

you're welcome, cya later and have a good night

OpenStudy (anonymous):

Thank you, You too